I'm trying to complete a fraction Magic Square. I must find out what "x" equals on each row and column of the square. I am given for the first row: x, -1/6 and 0; on the second row: -1/2, x, and 5/6; on the third row:1/3, x, and -1/3. I must find the fraction sum equaling the same number horizontally, vertically, and diagonally.
Answer by richwmiller(17219)

x , -1/6, 0
1/2, x , 5/6
1/3, x, -1/3
the third column adds up to 5/6-2/6=3/6=1/2
the third row is 0 with out x so x must be 1/2
1st row x-1/6+0=a, a-x=-1/6
2nd row 1/2+x+5/6=a,a-x=8/6
3rd row 1/3+x-1/3=a, a-x=0
They are not consistent.
There is no solution.